# Theory and Computation of Substructure Characteristic Modes

### v1

@@ Abstract
The problem of substructure characteristic modes is reformulated using a scattering matrix-based formulation, generalizing subregion characteristic mode decomposition to arbitrary computational tools. It is shown that the scattering formulation is identical to the classical formulation based on the background Green’s function for lossless systems. The scattering formulation, however, opens a variety of new subregion scenarios unavailable within previous formulations, including cases with lumped or wave ports or subregions in circuits. Thanks to its scattering nature, the formulation is solver-agnostic with the possibility to utilize an arbitrary full-wave method.
@@ I Introduction
Substructure characteristic mode decomposition involves dividing the scattering scenario into a controllable (or accessible, see [9, 10]) region and a background. In this approach, any structural modification, such as antenna design or selective excitation, is confined to the controllable region. An example is a patch antenna situated over a ground plane, where the ground plane has a significant impact on most characteristic modes, yet the designer can effectively influence only the patch design region [11, 9, 12]. The substructure characteristic mode decomposition focuses on modes most closely associated with the controllable region by using altered forms of operators describing the scattering problem. This makes substructure modes an attractive approach for studying the behavior of radiating devices affected by nearby objects like vehicles, electronic platforms, or biological tissues [13].
@@
Following the state-of-the-art procedures, substructure characteristic modes can be interpreted as a classical formulation of characteristic mode decomposition where the underlying Green’s dyadic includes the surroundings (background) of the studied region [14, 15, 16, 17]. One implementation of this approach using the method of moments (MoM) [18] formulations is based on a Schur complement [9, 19], which constructs a compressed impedance matrix for a scatterer in the presence of background objects.
@@
This paper expands the definition of substructure modes by introducing a scattering-based variant, formulated through a generalized eigenvalue problem of two scattering matrices [20, § 7.8.1], [21, § 4.3]. The first matrix accounts for the complete scattering problem, while the second represents the background. Following assumptions widely used in characteristic mode analysis, we assume losslessness, linearity, and time-invariance for all materials. The employed scattering matrices can be easily substituted with transition matrices (T-matrices) [22] or by scattering dyadics [20, § 4.3], which can be constructed by an arbitrary numerical technique [5, 6, 7]. The iterative algorithm [8] adapted here for the evaluation of substructure modes mitigates the computational burden associated with increasing electrical size and model complexity. This makes the proposed formulation applicable to various problems typically outside the scope of characteristic mode analysis, including microwave circuits and optical circuits spanning large ranges of electrical sizes with complex material distributions.
@@
The paper is organized as follows. Section II briefly recapitulates the meaning of substructure characteristic modes and the used nomenclature. Section III introduces a novel scattering-based formulation. Its equivalence to MoM-based substructure characteristic mode decomposition is shown in Section IV on cases when the formulation was known before, including perfect electric conductors, a hybrid of the method of moments and T-matrix method, and infinite ground plane. The advantages of scattering formulation are shown in Section V, dealing with challenging examples that are impractical to compute using previous approaches. The outcomes are discussed in Section VI, and the paper is concluded in Section VII.
@@ II Substructure Characteristic Modes
Insights into essential features of substructure characteristic modes can be obtained by comparing Fig. 1 and Fig. 2, where we adapt a PIFA-like example [9] made of perfect electric conductor (PEC), illustrating the characteristic modal significance of the entire region $\varOmega=\varOmega_{\mathrm{c}}\cup\varOmega_{\mathrm{b}}$ and substructure modal significance corresponding to the controllable region $\varOmega_{\mathrm{c}}$ in the presence of the background region $\varOmega_{\mathrm{b}}$.
@@
Figure 1: Modal significances $|t_{n}|$ computed for a PEC geometry adapted from [9]. The dimensions read $\ell=120\,\mathrm{mm}$, $w=60\,\mathrm{mm}$, $h=15\,\mathrm{mm}$, and $d=30\,\mathrm{mm}$. Impedance-based formulation [3] and scattering-based formulation (1) are employed with $\mathbf{S}_{\mathrm{b}}$ being the unit matrix.
Figure 2: Modal significances $|t_{n}|$ computed for the substructure $\varOmega_{\mathrm{c}}$ in the presence of the background scatterer $\varOmega_{\mathrm{b}}$. Dimensions are identical to Fig. 1. The impedance-based method relies on formulation [9], while the scattering-based formulation relies on (1).
@@ III Scattering Formulation of Substructure Characteristic Modes
Assume a controllable antenna region $\varOmega_{\mathrm{c}}$ and the surrounding (or background) region $\varOmega_{\mathrm{b}}$, as illustrated in Fig. 3. Let $\mathbf{S}_{\mathrm{b}}$ denote that scattering matrix [21, 20] for the background region $\varOmega_{\mathrm{b}}$, which connects incoming waves represented by coefficients collected in a vector $\mathbf{a}$ into outgoing waves represented by a vector $\mathbf{f}$. Similarly, let $\mathbf{S}$ denote the scattering matrix for the composite object $\varOmega_{\mathrm{c}}\bigcup\varOmega_{\mathrm{b}}$. The core hypothesis of this paper is that substructure characteristic modes are determined from the generalized eigenvalue problem
@@
where $\mathbf{a}_{n}$ are characteristic excitations and $s_{n}$ characteristic scattering eigenvalues. Throughout the remainder of the paper, we refer to these quantities as characteristic modes due to their equivalence with the characteristic currents $\mathbf{I}_{n}$ typically used in impedance-based formulations [3, 5]. This formulation generalizes the one defined by [24] and developed in [5] for objects in free space, which implicitly considers the background scattering matrix $\mathbf{S}_{\mathrm{b}}$ to be an identity matrix. It also covers the special case studied for periodic structures in [25], where the background scattering matrix $\mathbf{S}_{\mathrm{b}}$ is related to an implicit connection between plane waves on either side of a scattering surface, analogous to the S-parameters of a transmission line network. The scattering eigenvalues $s_{n}$ are related to modal significance $|t_{n}|$ and characteristic eigenvalues $\lambda_{n}$ as [5]
$t_{n}=\frac{s_{n}-1}{2}\quad\text{and }\lambda_{n}=-\operatorname{Im}\{t_{n}^{-1}\}=\mathrm{j}\frac{s_{n}+1}{s_{n}-1}.$ (2)
@@
which, for example, translates to the orthogonality of characteristic far fields and where ${}^{\mathrm{H}}$ denotes Hermitian transpose and $\delta_{mn}$ is the Kronecker delta. Equivalence between scattering and impedance formulations is supported by data shown in Figs. 1 and 2 comparing results from modal significances produced by (1) and an impedance-based formulation of substructure modes [9]. A mathematical derivation of this equivalence is outlined in Section IV and Appendix B.
@@ III-A Variations
where $\mathbf{f}_{n}=\mathbf{S}_{\mathrm{b}}\mathbf{a}_{n}=s_{n}^{\ast}\mathbf{S}\mathbf{a}_{n}$ represents the scattered field, $\mathbf{1}$ is the identity matrix, and ${}^{\ast}$ denotes complex conjugate. The two versions in (4) are equivalent and differ only in expressing the eigenvalue problem in incoming (excitation), $\mathbf{a}_{n}$, or outgoing (scattered or radiated), $\mathbf{f}_{n}$, waves. We note that the incident field $\mathbf{a}_{n}$ differ from the scattered field $\mathbf{f}_{n}$ except for the free-space case with $\mathbf{S}_{\mathrm{b}}=\mathbf{1}$ [5].
@@
which allows for the analysis of substructure characteristic modes solely in terms of transition matrices and characteristic number $t_{n}$. These formulations read
@@
$(2\mathbf{T}\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}+\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}+\mathbf{T})\mathbf{f}_{n}=t_{n}\mathbf{f}_{n}$ (7)
and does not follow a scheme analogous to the characteristic modes of isolated objects [5]. In contrast to (1), relations (4), (6), and (7) enable iterative matrix-free evaluation [8], an important factor when employing generic electromagnetic solvers for evaluating substructure characteristic modes for large problems. Details of the iterative solution is found in Appendix A. Moreover, relation (6) allows for interpreting the substructure scattered power as a power of the difference between the scattered field of the composite object $\mathbf{T}\mathbf{a}_{n}$ and the scattered field of the background $\mathbf{T}_{\mathrm{b}}\mathbf{a}_{n}$, i.e.
@@ IV Equivalence Between MoM and Scattering-Based Substructure Characteristic modes
The proposed formula (1) is equivalent to the impedance-based substructure modes [9], wherein a structure is separated into two (possibly connected) regions $\varOmega_{\mathrm{b}}$ (“background”) and $\varOmega_{\mathrm{c}}$ (“controllable”) and excitation is restricted to only the controllable region $\varOmega_{\mathrm{c}}$. A general scattering problem in the standard notation for electric field integral equation (EFIE)-formulations [18] reads
@@
where $\widetilde{\mathbf{Z}}=\widetilde{\mathbf{R}}+\mathrm{j}\widetilde{\mathbf{X}}=\mathbf{Z}_{\mathrm{c}\mathrm{c}}-\mathbf{Z}_{\mathrm{c}\mathrm{b}}\mathbf{Z}_{\mathrm{b}\mathrm{b}}^{-1}\mathbf{Z}_{\mathrm{b}\mathrm{c}}$ is an impedance matrix related to the problem-specific numerical Green’s function [14]. This numerical Green’s function captures the field-generating behavior of currents on the controllable region $\varOmega_{\mathrm{c}}$ in the presence of the “background” region $\varOmega_{\mathrm{b}}$.
@@
Appendix B then shows that matrix $\widetilde{\mathbf{T}}$ equals to the matrix used in (7). Similarly to full characteristic modes, the characteristic substructure current can be evaluated as [5]
$\mathbf{I}_{\mathrm{c}n}=t_{n}^{-1}\widetilde{\mathbf{Z}}^{-1}\widetilde{\mathbf{U}}_{1}^{\mathrm{H}}\mathbf{f}_{n}.$ (15)
IV-A Example — Demonstrating equivalence
@@
To demonstrate the numerical equivalence between the impedance and scattering formulations in (11) and (1), we employ a PEC structure reported in [9], which consists of a PIFA-like region connected to a finite ground plane, as depicted in the inset of Fig. 1. In Fig. 1, the characteristic modes of the full structure are computed without any separation into controllable and background regions, i.e., the entire region is controllable. Impedance and scattering operators are calculated using MoM, and the modal significances produced by the impedance and scattering formulations are numerically identical. Similarly, Fig. 2 displays modal significances when only part $\varOmega_{\mathrm{c}}$ of the structure is controllable (following [9]), and again the impedance (11) and scattering formulations (1) agree within numerical precision.
IV-B MoM & T-matrix hybrid

### v3

@@ Abstract
The problem of substructure characteristic modes is developed using a scattering matrix-based formulation, generalizing subregion characteristic mode decomposition to arbitrary computational tools. It is shown that the modes of the scattering formulation are identical to the modes of the classical formulation based on the background Green’s function for lossless systems under conditions where both formulations can be applied. The scattering formulation, however, opens a variety of new subregion scenarios unavailable within previous formulations, including cases with lumped or wave ports or subregions in circuits. Thanks to its scattering nature, the formulation is solver-agnostic with the possibility to utilize an arbitrary full-wave method.
@@ I Introduction
Substructure characteristic mode decomposition involves dividing the scattering scenario into a controllable (or accessible, see [9, 10]) design region and an uncontrollable background. In this approach, any structural modification, such as antenna design or selective excitation, is confined to the controllable region. An example is a patch antenna situated over a ground plane, where the ground plane has a significant impact on most characteristic modes, yet the designer can effectively influence only the patch  [11, 9, 12]. The substructure characteristic mode decomposition focuses on modes most closely associated with the controllable region by using altered forms of operators describing the scattering problem. This makes substructure modes an attractive approach for studying the behavior of radiating devices affected by nearby objects like vehicles, electronic platforms, or biological tissues [13].
@@
Following the state-of-the-art procedures, substructure characteristic modes can be interpreted as a classical formulation of characteristic mode decomposition where the underlying Green’s dyadic includes the surroundings (background) of the studied region [14, 15, 16, 17]. For a few special cases, the Green’s dyadic can be computed analytically (e.g., infinite ground planes), however most practical problems require numerical solutions. One implementation of this approach using the method of moments (MoM) [18] formulations is based on a Schur complement [9, 19], which constructs a compressed impedance matrix for a scatterer in the presence of background objects. Several variations of this technique have been developed to account for a variety of specific background problems, all following the Schur complement method applied to various integral equation formulations [20, 21, 22, 23, 24, 25, 26, 27, 28].
@@
This paper proposes an alternate approach to the computation of substructure modes for lossless problems without the explicit requirement of integral equations by introducing a scattering-based variant, formulated through a generalized eigenvalue problem of two scattering matrices [29, § 7.8.1], [30, § 4.3]. The first matrix accounts for the complete scattering problem, while the second represents the background. Unlike previous methods, this approach is general and opens up the analysis of previously inaccessible scenarios such as bi-anisotropic medium and ports (see Sec. VI).The employed scattering matrices can be easily substituted with transition matrices (T-matrices) [31] or by scattering dyadics [29, § 4.3], which can be constructed by an arbitrary numerical technique [5, 6, 7]. Furthermore, the proposed algorithm can be accelerated through the use of an adapted form of a previously-reported iterative algorithm [8], which mitigates the computational burden associated with increasing electrical size and model complexity in substructure problems.
The scattering-based and MoM-based substructure formulations are shown to be identical in the sense that they yield the same modes in cases where both methods are applicable. Nevertheless, there are cases where the MoM-based variant is not easily applicable or not advantageous to use. This makes the proposed formulation applicable to various problems typically outside the scope of characteristic mode analysis, including microwave circuits and optical circuits spanning large ranges of electrical sizes with complex material distributions. In summary, the proposed method inherits all of the properties of the scattering formulation of characteristic modes [2, 5] including the most impactful generalization that if the scattering response of an object and its background can be computed (by any means), then substructure characteristic modes can be computed, regardless of the background or object’s complexity and whether or not integral equations were used.
@@
The paper is organized as follows. Section II briefly reviews the basic concept of substructure characteristic modes. Section III introduces the scattering-based formulation and its equivalence to MoM-based substructure characteristic mode decomposition is shown in Section IV. Several examples are presented, with emphasis on those which can be computed both by previous methods and the proposed scattering-based technique. These include a simple object made of perfect electric conductor (PEC), a hybrid of the method of moments and T-matrix method, and infinite ground plane. Computationally challenging examples are shown in Section V. The outcomes are discussed in Section VI, and the paper is concluded in Section VII.
@@ II Substructure Characteristic Modes
Insights into essential features of substructure characteristic modes can be obtained by comparing Fig. 1 and Fig. 2, where we adapt a PIFA-like example [9] made of PEC [^1] We study variations of this example throughout the remainder of the manuscript using its original dimensions and frequency analysis range from [9], though the absence of losses means that all problems can be scaled in physical size to cover arbitrary frequency ranges without alteration to the theoretical technique or numerical results., illustrating the characteristic modal significance of the entire region $\varOmega=\varOmega_{\mathrm{c}}\cup\varOmega_{\mathrm{b}}$ and substructure modal significance corresponding to the controllable region $\varOmega_{\mathrm{c}}$ in the presence of the background region $\varOmega_{\mathrm{b}}$.
@@
Figure 1: Modal significances $|t_{n}|$ computed for a PEC geometry adapted from [9] using the scattering-based formulation (1) with $\mathbf{S}_{\mathrm{b}}$ being unit matrix, i.e., $\mathbf{S}_{\mathrm{b}}=\mathbf{1}$, (solid lines), and the impedance-based formulation (dots) [3]. The dimensions read $\ell=120\,\mathrm{mm}$, $w=60\,\mathrm{mm}$, $h=15\,\mathrm{mm}$, and $d=30\,\mathrm{mm}$.
Figure 2: Modal significances $|t_{n}|$ computed for the substructure $\varOmega_{\mathrm{c}}$ in the presence of the background scatterer $\varOmega_{\mathrm{b}}$. Both regions are PEC with dimensions identical to the structure in Fig. 1. The impedance-based method relies on formulation [9], while the scattering-based formulation relies on (1).
@@ III Scattering Formulation of Substructure Characteristic Modes
Assume a controllable antenna region $\varOmega_{\mathrm{c}}$ and the surrounding (or background) region $\varOmega_{\mathrm{b}}$, as illustrated in Fig. 3. Let $\mathbf{S}_{\mathrm{b}}$ denote that scattering matrix [30, 29] for the background region $\varOmega_{\mathrm{b}}$, which connects incoming waves represented by coefficients collected in a vector $\mathbf{a}$ into outgoing waves represented by a vector $\mathbf{f}$. Similarly, let $\mathbf{S}$ denote the scattering matrix for the composite object $\varOmega_{\mathrm{c}}\cup\varOmega_{\mathrm{b}}$. These scattering matrices are general, and the basis functions used to expand the incident and reflected waves depend on the specific problem. For a free-space scattering problem it is convenient, though not necessary, to use spherical vector waves as basis functions [29]. In this case, the column vectors $\mathbf{f}$ and $\mathbf{a}$ contain the spherical wave expansion coefficients, similar to a multipole expansion.
The core hypothesis of this paper is that substructure characteristic modes are determined from the generalized eigenvalue problem [33]
@@
where $\mathbf{a}_{n}$ are characteristic excitations and $s_{n}$ characteristic scattering eigenvalues. Characteristic scattered fields are represented by vectors $\mathbf{f}_{n}=\mathbf{S}_{\mathrm{b}}\mathbf{a}_{n}$ which collect expansion coefficients of the scattered field in the same basis.
The dimension of the square matrices $\mathbf{S}$ and $\mathbf{S}_{\mathrm{b}}$ are the same, with conservative estimates available for the number of entries based on the electrical size of the problems being studied [^2] If the background structure is much smaller than that of the controllable region, then a smaller dimension can be used to accelerate computation of the background scattering matrix $\mathbf{S}_{\mathrm{b}}$ before augmenting it with an identity matrix to match the dimension of matrix $\mathbf{S}$ which represents the complete structure., such as the electrical radius for expansions in spherical waves [34, 35]. This formulation generalizes the one defined by [36] and developed in [5] for objects in free space, which implicitly considers the background scattering matrix $\mathbf{S}_{\mathrm{b}}$ to be an identity matrix.
The formulation above also covers the special case studied for periodic structures in [37], where the background scattering matrix $\mathbf{S}_{\mathrm{b}}$ is related to an implicit connection between plane waves (Floquet modes) on either side of a scattering surface, analogous to the S-parameters of a transmission line network. The background problem for periodic systems can also be modified to include a ground plane or supporting material layers [37] and the use of periodic boundary conditions leads to much more restrictive guidelines on the appropriate dimension of the scattering matrices based on the number of propagating Floquet modes [38].
The eigenvalues $s_{n}$ are related to modal significances $|t_{n}|$ and characteristic eigenvalues $\lambda_{n}$ as [5]
$t_{n}=\frac{s_{n}-1}{2}\quad\text{and }\lambda_{n}=\mathrm{j}(1+t_{n}^{-1})=\mathrm{j}\frac{s_{n}+1}{s_{n}-1}.$ (2)
@@
which, for example, translates to the orthogonality of characteristic far fields and where H denotes Hermitian transpose and $\delta_{mn}$ is the Kronecker delta. Equivalence between the modes from the scattering and impedance formulations is supported by data shown in Figs. 1 and 2 comparing results from modal significances produced by (1) and an impedance-based formulation of substructure modes [9]. A mathematical derivation of this equivalence is outlined in Section IV and Appendix B.
Throughout the remainder of the paper, we use the quantities $\mathbf{a}_{n}$ and $s_{n}$ as representations of characteristic modes due to their direct relation to the characteristic currents $\mathbf{I}_{n}$ typically used in impedance-based formulations [3, 5]. These characteristic currents are typically referred to as ”excitation independent”, but they can always be driven by an appropriate excitation $\mathbf{V}_{n}=\mathbf{Z}\mathbf{I}_{n}$ (obtained using the impedance matrix $\mathbf{Z}$) or, equivalently, the superposition of spherical waves described by the vector $\mathbf{a}_{n}$. Details on recovering the characteristic currents are given in Section IV-A. Other characteristic quantities, such as fields, are computed by exciting the system with an excitation described by $\mathbf{a}_{n}$ and measuring the quantity of interest, using whatever solver was used to construct the scattering matrices $\mathbf{S}$ and $\mathbf{S}_{\mathrm{b}}$. The explicit mapping between incident fields and modal currents within MoM-based solvers is discussed in Section IV.
@@ III-A Variations
where $\mathbf{f}_{n}=\mathbf{S}_{\mathrm{b}}\mathbf{a}_{n}=s_{n}^{\ast}\mathbf{S}\mathbf{a}_{n}$ represents the scattered field, $\mathbf{1}$ is the identity matrix, and  denotes complex conjugate. The two versions in (4) are equivalent and differ only in expressing the eigenvalue problem in incoming (excitation), $\mathbf{a}_{n}$, or outgoing (scattered or radiated), $\mathbf{f}_{n}$, waves. We note that the incident field $\mathbf{a}_{n}$ differ from the scattered field $\mathbf{f}_{n}$ except for the free-space case with $\mathbf{S}_{\mathrm{b}}=\mathbf{1}$ [5].
@@
which allows for the analysis of substructure characteristic modes solely in terms of transition matrices and characteristic numbers $t_{n}$. These formulations read
@@
$(2\mathbf{T}\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}+\mathbf{T}_{\mathrm{b}}^{\mathrm{H}}+\mathbf{T})\mathbf{f}_{n}=t_{n}\mathbf{f}_{n},$ (7)
which involve products of transition matrices, unlike the characteristic mode formulation of isolated objects [5]. In contrast to (1), relations (4), (6), and (7) enable iterative matrix-free evaluation [8], an important factor when employing generic electromagnetic solvers for evaluating substructure characteristic modes for large problems. Details of the iterative solution are found in Appendix A. Moreover, relation (6) allows for interpreting the substructure scattered power as a power of the difference between the scattered field of the composite object $\mathbf{T}\mathbf{a}_{n}$ and the scattered field of the background $\mathbf{T}_{\mathrm{b}}\mathbf{a}_{n}$, i.e.
@@ IV Equivalence Between MoM and Scattering-Based Substructure Characteristic modes
The proposed formula (1) yields the same modes as the impedance-based substructure formulation [9], wherein a structure is separated into two (possibly connected) regions $\varOmega_{\mathrm{b}}$ and $\varOmega_{\mathrm{c}}$, and excitation is restricted to only the controllable region $\varOmega_{\mathrm{c}}$ [33]. A general scattering problem in the standard matrix representation of surface and / or volumetric electric field integral equation (EFIE)-formulations [18] reads
@@
where $\widetilde{\mathbf{Z}}=\widetilde{\mathbf{R}}+\mathrm{j}\widetilde{\mathbf{X}}=\mathbf{Z}_{\mathrm{c}\mathrm{c}}-\mathbf{Z}_{\mathrm{c}\mathrm{b}}\mathbf{Z}_{\mathrm{b}\mathrm{b}}^{-1}\mathbf{Z}_{\mathrm{b}\mathrm{c}}$ is an impedance matrix related to the problem-specific numerical Green’s function [14]. This numerical Green’s function captures the field-generating behavior of currents on the controllable region $\varOmega_{\mathrm{c}}$ in the presence of the background region $\varOmega_{\mathrm{b}}$.
@@
Appendix B then shows that matrix $\widetilde{\mathbf{T}}$ equals to the matrix used in (7).
The characteristic current can be obtained from the characteristic excitation as
$\mathbf{I}_{n}=\begin{bmatrix}\mathbf{I}_{\mathrm{b}n}\\ \mathbf{I}_{\mathrm{c}n}\end{bmatrix}=\mathbf{Z}^{-1}\mathbf{U}_{1}^{\mathrm{H}}\mathbf{a}_{n}-\begin{bmatrix}\mathbf{Z}_{\mathrm{b}\mathrm{b}}^{-1}\mathbf{U}_{1\mathrm{b}}^{\mathrm{H}}\\ \mathbf{0}\end{bmatrix}\mathbf{a}_{n},$ (15)
i.e., by subtracting the current induced by excitation $\mathbf{a}_{n}$ on the background object from the current induced on the composite object. Employing the abbreviations introduced in this section, relation (15) is identical to
$\mathbf{I}_{\mathrm{c}n}=\dfrac{1}{t_{n}}\widetilde{\mathbf{Z}}^{-1}\widetilde{\mathbf{U}}_{1}^{\mathrm{H}}\mathbf{f}_{n}.$ (16)
IV-A Example — Equivalence and Substructure Interpretation
@@
To demonstrate the numerical equivalence between the modes from the impedance and scattering formulations in (11) and (1), we employ a PEC structure reported in [9], which consists of a PIFA-like region connected to a finite ground plane, as depicted in the inset of Fig. 1. In Fig. 1, the characteristic modes of the full structure are computed without any separation into controllable and background regions, i.e., the entire region is controllable. Impedance and scattering operators are calculated using MoM, and the modal significances produced by the impedance and scattering formulations are numerically identical. Similarly, Fig. 2 displays modal significances when only part $\varOmega_{\mathrm{c}}$ of the structure is controllable (following [9]), and again the impedance (11) and scattering formulations (1) agree within numerical precision.
Figure 4: Comparison between modal significances $|t_{n}|$ for the PIFA structure in Fig. 1 (solid) and a ground plane (dashed). No background is considered, i.e., $\mathbf{S}_{\mathrm{b}}=\mathbf{1}$. The colors are paired according to the similarity in radiation diagrams. When evaluating substructure characteristic modes in Fig. 2, the ground plane (dashed lines) is taken as a background. All dimensions are the same as in Fig. 1.
The transformation of full-structure modes, shown in Fig. 4, into substructure modes in Fig. 2 can be intuitively explained using the excitation and radiated fields described in (8).
To illustrate this, consider the resonances around $1\mathrm{\,GHz}$ for both full structures in Fig. 4. These resonances correspond to the fundamental mode of each structure where the far-field closely resembles that of an electric dipole, with characteristic currents flowing along the longest dimension. The resonant frequency of the full PIFA structure is slightly lower than that of the ground plane alone, mainly due to the increased effective length of the overall configuration.
Assuming similar characteristic excitation $\mathbf{a}_{n}$ for both modes near resonance, (8) suggests that substructure characteristic modes result from the difference in scattering – both amplitude and phase – between the full PIFA structure and the ground plane “background”. This difference is reflected in the substructure eigenvalues in Fig. 2, which is given from the differences of the eigenvalues in Fig. 4. Consequently, no resonance is observed near $1\mathrm{\,GHz}$ in Fig. 2; however, the dominant substructure mode exhibits a peak with a modal significance of $|t_{n}|\approx 0.6$. This behavior is attributed to the frequency shift between the black lines in Fig. 4 and the resulting imperfect cancellation of eigenvalues.
The second characteristic mode, shown by the blue traces in Fig. 4, similarly has a far field characterized by that of an electric dipole, with characteristic currents flowing along the width dimension. The characteristic traces are nearly identical for both structures, resulting in a significant reduction for the substructure modal significance, as indicated by (8). Finally, the resonance of the full structure in Fig. 4 around $1.6\mathrm{\,GHz}$ is not supported by the ground plane (see the full red line in Fig. 4) and is therefore replicated in the substructure data in Fig. 2.
These simple interpretations for the substructure characteristic modes rely on the assumption that the excitations $\mathbf{a}_{n}$ are approximately the same for the entire structure and the background structure. This approximation holds well over large bandwidths for electrically small structures but only for narrow bandwidths for electrically large structures.
IV-B Computational complexity
In contrast to the MoM matrix formulation (11), the scattering-based substructure formulation (1) is solver-agnostic. The scattering formulation can thus take advantage of the rapid advances in computational electromagnetics (CEM) to solve scattering problems and, hence, effectively evaluate characteristic modes. A theoretical comparison of the computational complexity between the classical MoM formulation and the scattering-based formulation is presented for two idealized cases in Table I.
Table I: Comparison between memory usage and number of operations for the MoM formulation (10) together with (11) and iterative matrix-free MoM formulations [39, 40] for the scattering formulation (1).
MoM Scattering
[Uncaptioned image] mem. $(k\ell)^{4}$ $(k\ell)^{2}$
ops. $(k\ell)^{6}$ $N_{\mathrm{cm}}N_{\mathrm{it}}(k\ell)^{2}\log(k\ell)$
[Uncaptioned image] mem. $(k\ell)^{6}$ $(k\ell)^{3}$
ops. $(k\ell)^{9}$ $N_{\mathrm{cm}}N_{\mathrm{it}}(k\ell)^{3}\log(k\ell)$
The background object in the first case consists of a PEC plate with side lengths $\ell$, while the second case features a dielectric cube with side lengths $\ell$ and a spatially dependent relative permittivity $\varepsilon_{\mathrm{r}}(\boldsymbol{r})$. For simplicity, the controllable region is assumed to be a square PEC plate with side lengths $\ell_{\mathrm{c}}\ll\ell$. Assuming a discretization using 10 points per wavelength results in approximately $N_{\mathrm{b}}\approx 10\ell_{\mathrm{b}}/\lambda\approx 1.6k\ell$ free-space discretization points per side. The memory usage of the MoM for the PEC plate is $N_{\mathrm{b}}^{4}\sim(k\ell)^{4}$ matrix elements for $\mathbf{Z}_{\mathrm{b}\mathrm{b}}$. The computational complexity for evaluation of the Schur complement (11) is proportional to $(k\ell)^{6}$. For the dielectric cube, the memory and complexity are proportional to $(k\ell)^{6}$ and $(k\ell)^{9}$, respectively, assuming the use of a volumetric electric field integral equation.
In contrast, the scattering problems underlying (1) can be solved using arbitrary solvers. Taking advantage of matrix-free MoM formulations, such as the Multi-Level Fast Multipole Method (MLFMM), reduces the memory requirement to $(k\ell)^{2}$ [39] for the PEC case. The characteristic modes are solved iteratively as outlined in Appendix A, giving a computational complexity proportional to the number of evaluated characteristic modes $N_{\mathrm{cm}}$, MLFMM iterations $N_{\mathrm{it}}$, and $(k\ell)^{2}\log(k\ell)$. The corresponding heterogeneous dielectric cube can, for instance, be solved using FFT-based volumetric MoM [40], with a memory requirement of $(k\ell)^{3}$ and similar computational complexity.
Both cases in Table I demonstrate the unfavorable memory and computational complexity of MoM systems (10) for electrically large structures. The memory scaling of $(k\ell)^{6}$ for volumetric objects makes the MoM system (10) impractical for objects larger than a few wavelengths. In contrast, the full range of modern CEM techniques (including MoM, FEM, and FDTD) is directly applicable to the scattering-based formulation (1). This example illustrates some advantages of (1) over the classical approach (11). However, it is important to note that we do not claim that the scattering-based approach is universally superior. There are numerous situations where the classical method is computationally efficient and may be preferable, such as for electrically small controllable and background objects represented by a low number of basis functions (e.g., the example shown in Fig. 2). The scaling in Table I is based on the use of a Schur complement (11) for the MoM matrices. However, recent advancements can potentially improve the computational efficiency of the impedance-based substructure characteristic modes [41], similar to how the MLFMM is applied to full characteristic modes in [42].
IV-C MoM and T-matrix Hybrid